# No Free Lunch Theorem for Privacy-Preserving LLM Inference

### v1

@@ 1 Introduction
Although LLMs bring great benefits, they have also raised privacy concerns. When users interact with LLMs, their queries, which may encompass confidential personal or company information, are exposed to the host of LLMs. If users’ queries are not adequately secured and LLM vendors are not effectively regulated, there is a risk that user data could be misused or accessed by unauthorized individuals, leading to identity theft or other privacy violations. This is especially worrying considering the sensitivity of the information, such as financial details or health-related discussions, shared with AI chatbots or Agents.
Therefore, preserving privacy is of great importance when prompting LLMs for inference.
In this study, our focus lies on a typical LLM inference setting, where the LLM is black-box, and its architecture, model parameters, and inference details are concealed within commercial APIs and user interfaces. In light of this, a natural approach to protecting the privacy of prompts is to add randomization to prompts. Several randomization techniques [4, 5, 6] have been proposed to address this concern and have shown empirical effectiveness in mitigating privacy leakage resulting from exposed prompts. However, these approaches inevitably incur a certain degree of utility loss compared to prompting LMs without any protection. Moreover, existing research lacks the theoretical analysis necessary to quantify the compromise of the reduction in utility and exposure of privacy. The tension between reducing the exposure of private information and the detriment to utility propels our exploration to investigate the research inquiry: "Theoretically, is there possibility of developing some protective methods that can achieve minimal exposure of private information and detriment to utility at the same time when prompting an LLM?" Our primary finding is encapsulated in the Theorem 4.4 within context of our proposed framework for privacy-preserving LM inference, presenting a counterargument to this question. We named it as No-Free-Lunch Theorem.
@@ 2 Related Work
During the LLM inference, both the client and the LLM server have the potential to act as adversaries, compromising each other’s privacy [8]. In this study, we focus on a scenario where the LLM server is the adversary and can infer the client’s private information by analyzing the prompt provided by the client (see Section 3.1). In this section, we provide a concise overview of existing research concerning attack of privacy, safeguarding measures, and the balance of preserving utility against protecting privacy within the realm of large language model (LLM) inference.
@@ 3.4 Utility Loss and Privacy Leakage
$\epsilon_{p}=R(\breve{P})-R(\widetilde{P})$ (1)
@@ 4 No Free Lunch Theorem for Privacy-Preserving LLM Inference
Lemma 4.1 (Theoretical Relationship between Privacy Leakage and Distortion Extent).
@@
Let $\epsilon_{p}$ be defined in Eq. (2). Let $\widetilde{P}$ and $\breve{P}$ represent prompt embedding’s distribution before and after distorted. Given the preceding conditions, it follows
@@ 5 Conclusion
To answer this question, we propose a privacy-preserving LLM inference framework. Within this framework, we propose a No-Free-Lunch Theorem for privacy-preserving LLM inference, demonstrating that the weighted summation of these factors exceeds a constant which is contingent on specific problem and non-zero. This highlights inevitable loss of utility when leakage budget of privacy is excessively constrained.
@@ 6 ACKNOWLEDGMENTS
This work was supported by the National Science and Technology Major Project under Grant 2022ZD0115301.
@@
Lemma A.5 (Theoretical Relationship between Privacy Leakage and Distortion Extent).
@@
Let $\epsilon_{p}$ be defined in Eq. (2). Let ${\widetilde{P}}$ and $\breve{P}$ represent the distribution of the embedding before and after being protected. Then we have,

### v3

@@ 1 Introduction
Figure 1: Privacy Leakage in Prompts. In Case 1, the prompt contains sensitive user information, including: name (John Smith), address (1234 Elm Street, Springfield), and medical condition (high blood pressure). This enables the LLM to identify and extract private data, potentially leading to privacy leakage. Such information risks exposing the user’s identity, location, and health details, which could be misused for identity theft, targeted advertising, or discrimination. In contrast, the prompt in Case 2 excludes sensitive information, significantly reducing the risk of privacy leakage. This comparison highlights how sensitive information in user prompts can be easily identified by LLMs.
While LLMs offer significant benefits, they also raise important privacy concerns. User prompts, which are central to interactions with LLMs, often include sensitive information such as personal identity, background details, and other confidential data, as illustrated in Figure 1. When users do not employ privacy protection measures, any prompt containing sensitive information can be easily recognized and extracted by the server. Even with certain privacy protection techniques in place, servers may still deploy various attack methods—such as embedding inversion or model inference—to infer sensitive user data. These threats have been substantiated in numerous studies (see Section 2.1). However, overly stringent privacy protections can compromise the utility of LLMs, while insufficient protection leaves user privacy vulnerable. Therefore, our research aims to strike a balance by ensuring privacy protection for user prompts during the LLM inference phase, with minimal impact on model utility.
Therefore, preserving privacy is crucial when prompting LLMs for inference. In this context, a natural approach to safeguarding the privacy of prompts is to introduce randomization.
This study focuses on a typical LLM inference setting where the LLM operates as a black box, concealing its architecture, model parameters, and inference details within commercial APIs and user interfaces. To address this issue, several randomization techniques [4, 5, 6] have been proposed, demonstrating empirical effectiveness in mitigating privacy leakage resulting from exposed prompts. However, these approaches inevitably incur a certain degree of utility loss compared to prompting LMs without any protection. Moreover, existing research lacks the theoretical analysis necessary to quantify the compromise of the reduction in utility and exposure of privacy. The tension between reducing the exposure of private information and the detriment to utility propels our exploration to investigate the research inquiry: "Theoretically, is there possibility of developing some protective methods that can achieve minimal exposure of private information and detriment to utility at the same time when prompting an LLM?" Our primary finding is encapsulated in the Theorem 4.4 within context of our proposed framework for privacy-preserving LM inference, presenting a counterargument to this question. We named it as No-Free-Lunch Theorem.
@@ 2 Related Work
During the LLM inference, both the client and the LLM server have the potential to act as adversaries, compromising each other’s privacy [8, 9]. In this study, we focus on a scenario where the LLM server is the adversary and can infer the client’s private information by analyzing the prompt provided by the client (see Section 3.1). In this section, we provide a concise overview of existing research concerning attack of privacy, safeguarding measures, and the balance of preserving utility against protecting privacy within the realm of large language model (LLM) inference.
@@ 3.1 Threat Model
Additionally, while our primary focus is on scenarios where the LLM server acts as the adversary, we also consider the potential risks posed by malicious clients. Malicious clients may attempt to infer other clients’ private data by exploiting shared inference results, model updates, or interactions within collaborative learning frameworks. Our framework is designed to mitigate such risks by implementing a robust isolation mechanism, where each client’s prompt undergoes random perturbation before being processed by the LLM. This approach ensures that the LLM server, acting as an intermediary, does not directly access unprotected client data, thereby preventing malicious clients from inferring sensitive information about other users. As a result, our framework is adaptable to scenarios involving both a potentially adversarial server and malicious clients, ensuring comprehensive protection of client privacy.
@@ 3.4 Utility Loss and Privacy Leakage
$\epsilon_{p}=R(\widetilde{P})-R(\breve{P})$ (1)
@@ 4 No Free Lunch Theorem for Privacy-Preserving LLM Inference
Lemma 4.1 (Theoretical Relationship between Recovery Extent and Distortion Extent).
@@
Let $\epsilon_{p}$ be defined in Definition 3.1. Let $\widetilde{P}$ and $\breve{P}$ denote the distorted prompt embedding distribution and the independent random prompt embedding distribution. Given the preceding conditions, it follows
@@
5 Experiment
In this section, we first introduce InferDPT, a privacy-preserving algorithm for LLM inference proposed by [5], which safeguards user prompt privacy by injecting controlled noise into the embedding vectors of the original prompts, with the aim of exploring the inherent trade-off between privacy leakage and utility loss. Next, we outline the experimental setup and apply our definitions of privacy leakage and utility loss to evaluate how these factors evolve as the level of privacy protection is reduced.
5.1 InferDPT Algorithm Overview
The InferDPT framework is an innovative application of differential privacy[32] in LLM inference, designed to enable privacy-preserving text generation, particularly for remote black-box LLMs. The framework comprises two main components: a perturbation module and an extraction module.
Perturbation Module: The perturbation module protects privacy by introducing noise to the embedding vectors of each token in the original document, leveraging a differential privacy mechanism. The noise level is controlled by the privacy budget, $\epsilon$, with smaller values of $\epsilon$ providing stronger privacy guarantees. Additionally, the module generates a random adjacency list for each token, selecting replacement tokens to further strengthen privacy protection. This approach maintains the semantic coherence of the text while ensuring user privacy. The final perturbed document is created by applying these steps to each token in the original text. Subsequently, the user combines the perturbed document with task instructions to form a complete prompt, which is submitted to the remote black-box LLM. The LLM then generates a privacy-preserving output based on the perturbed input.
Extraction Module: The extraction module employs a local LLM, which is lightweight and deployable, while less capable than the remote black-box LLM. This local LLM generates text based on the original document and the privacy-preserving output as a reference.
The InferDPT algorithm offers several key advantages, including effective user prompt privacy protection, resistance to embedding inversion attacks, and its high-quality text generation. Furthermore, the algorithm is highly adaptable to various local LLM and deployment environments. Its modular design—comprising a perturbation module and an extraction module—enhances scalability, accommodating diverse privacy requirements and application scenarios. Specifically, InferDPT enables differential privacy across a wide range of text data scales and complexities. InferDPT incurs lower computational overhead compared to homomorphic encryption approaches (e.g., CipherGPT[24]), effectively balancing computational costs through dynamic sampling. Experimental results show that InferDPT maintains text generation quality on par with non-private GPT-4 across multiple datasets and a range of $\epsilon$ values.
However, the InferDPT algorithm also has some limitations. The quality of the generated text depends on the choice of the local LLM and the design of prompt. More importantly, it requires a trade-off between privacy protection and text generation quality, especially when strong privacy guarantees are applied. Additionally, deploying InferDPT requires certain hardware resources to run the local LLM, which may not be user-friendly for ordinary users.
Based on the above analysis, we believe that InferDPT is highly suitable for validating our theory. We will build upon this algorithm to explore the trade-off between privacy leakage and utility loss.
5.2 Experiment Setup
Dataset: We used the CNN/Daily Mail dataset[33] to perform an open-ended text generation task. Following previous works[5][34][35], perturbed document containing 50 tokens is provided to the remote black-box LLM or the local LLM, with additional task descriptions appended to form a complete prompt (all prompts can be found in 5.3). LLM then generates the following 100 tokens based on this prompt.
Model: We used GPT-3.5-turbo as the remote black-box LLM and Vicuna-7b-4bit as the local LLM. Both models were configured with a temperature setting of $0.5$ and a maximum token generation limit (max_tokens) of $150$.
Quantifying Privacy Leakage $\epsilon_{p}$: We first randomly select 50 tokens from the vocabulary to simulate the remote black-box LLM’s attempt to recover the perturbed document through random guessing. However, we find that directly calculating the recovery extent $R$ as defined in Definition 3.1 is problematic because the black-box LLM’s responses do not have a direct correspondence with the perturbation tokens. To address this issue, we use cosine similarity as a measure of the recovery extent. Cosine similarity is defined as $cos(\theta)=\frac{A\cdot B}{|A||B|}$, which is consistent with the definition of recovery extent in Definition 3.1, where $cos(\theta)$ closer to 0 indicates a higher recovery extent, and $cos(\theta)$ closer to 1 indicates a lower recovery extent.
We calculate the cosine similarity $cos(\theta))_{1}$ between the randomly guessed document and the original document, denoted as $R(\breve{P})$, which represents the recovery extent under random guessing. Then, we send the perturbed document to the remote black-box LLM, which attempts to recover the original document, and calculate the cosine similarity $cos(\theta))_{2}$ between the recovered document and the original document, denoted as $R(\widetilde{P})$, which represents the recovery extent of the remote black-box LLM. According to Definition 3.1, we can obtain $\epsilon_{p}$. Our expected result is that as the privacy budget $\epsilon$ increases, $R(\breve{P})$ remains stable, while $R(\widetilde{P})$ gradually increases, indicating that the remote black-box LLM’s ability to recover the user’s privacy increases, leading to an increase in privacy leakage $\epsilon_{p}$.
Quantifying Utility Loss $\epsilon_{u}$: According to our definition of utility loss in 3.2, Utility loss quantifies the difference in performance between LLM inference with and without privacy protection. Specifically, we first send the original document to the remote black-box LLM for a text continuation task, obtaining 100 generated tokens. We calculate the utility of this output as $U(P)$. Next, we perturb the original document to create a protected version and send it to the remote black-box LLM again, generating another set of 100 tokens. These 100 tokens, along with the original prompt, are then fed into the local LLM to extract information and calculate its utility metrics, denoted as $U(\widetilde{P})$. The utility loss of InferDPT is quantified as the difference between these two utilities.
To evaluate utility, we employ several metrics commonly used in open-ended text generation tasks, including BERTScore [36], BLEU, Keyword Coverage, Semantic Similarity, Diversity, Coherence, and ROUGE (ROUGE-1, ROUGE-2, and ROUGE-L). These metrics are described in detail as follows:
  • BERTScore: Measures the semantic similarity between the generated and reference texts using BERT embeddings, reflecting content consistency. It is particularly effective for evaluating semantic fidelity, as it captures the contextual meaning of words through pre-trained BERT models.
  • BLEU: Evaluates the precision of n-grams in the generated text compared to the reference text, providing a measure of fluency and faithfulness. BLEU is widely used in tasks such as machine translation and text generation.
  • Keyword Coverage: Assesses the extent to which the generated text retains key concepts or terms from the reference text. This ensures that essential information or keywords are accurately represented in the output.
  • Semantic Similarity: Beyond BERTScore, this metric captures deeper contextual and conceptual alignment between the generated and reference texts, providing a complementary perspective on semantic fidelity.
  • Diversity: Measures the variety and richness of the generated text, ensuring that the model avoids repetitive outputs. This is particularly important under privacy constraints, which can sometimes reduce variability in the content.
  • Coherence: Evaluates the logical flow and structural consistency of the generated text, ensuring that the output aligns with the expected narrative or topic and maintains readability.
  • ROUGE Scores (ROUGE-1, ROUGE-2, and ROUGE-L): Measure the overlap of n-grams between the generated and reference texts. ROUGE-1 focuses on unigram overlap, ROUGE-2 on bigrams, and ROUGE-L considers the longest common subsequences. These metrics are widely used for assessing textual accuracy, coverage, and relevance.
We divided privacy protection into 24 distinct levels and calculated both privacy leakage and utility loss for each level. By plotting privacy leakage on the x-axis and utility loss on the y-axis, we mapped points corresponding to different privacy budgets or noise intensities. As expected, utility loss decreases gradually as privacy leakage increases, indicating a reduction in privacy protection strength.
5.3 Design of Prompts for the InferDPT Framework
In our study, we employed different quantification strategies for measuring privacy leakage and utility loss, which required distinct prompt designs. Specifically, for quantifying privacy leakage, we tasked the remote black-box LLM with attempting to recover the perturbed documents. For quantifying utility loss, we instructed the remote black-box LLM to complete continuation tasks based on the input text, and then used the local LLM to extract relevant content from the original document.
5.3.1 Prompts for Quantifying Privacy Leakage
We designed the following prompt to instruct the remote black-box LLM to act as an attacker attempting to recover the perturbed documents:
You are a text recovery expert. Your task is to recover the original tokens that were replaced by their semantically similar alternatives.
Examples:
  • Original: The company announced record profits in the third quarter.
    Perturbed: Business speak happy money three time part.
    Recovered: The business reported positive financial results in the third period.
  • Original: Scientists discover new species in Amazon rainforest.
    Perturbed: People find thing green tree place.
    Recovered: Researchers identified a new organism in the tropical forest.
Instructions:
  • Perturbed:
  • You should only respond with the recovered tokens.
  • Output format: [token1] [token2] [token3] …
5.3.2 Prompts for Quantifying Utility Loss
To quantify utility loss, we first provide the original document to the remote black-box LLM to generate a continuation. Next, the remote black-box LLM is tasked with generating a continuation for the perturbed document. To evaluate the utility loss, we design prompts for the local LLM to extract the continuation content from the remote black-box LLM’s responses, using the original document as a reference. Accordingly, this section presents the prompts designed for both the remote black-box LLM’s continuation task and the local LLM’s extraction task.
Prompt for Remote Black-box LLM:
You are a creative writer skilled in continuing stories and texts in a natural and engaging way.Your task is to generate a coherent continuation that matches the style and context of the original text.
Continue this text in a natural and engaging way, maintaining the same style and tone. Write 2-3 sentences that flow naturally from the original text.
——Original Text:
——[your continuation]
Prompt for Local LLM:
Analyze the coherence between the "Prefix Text" and "Perturbed Generation", then follow these rules:
1. If the "Perturbed Generation" meets ALL these criteria:
- Continues the story/topic from the "Prefix Text" naturally
- Maintains consistent facts and details
- Uses similar tone and style
→ Output the "Perturbed Generation" exactly as is
2. Otherwise, if ANY criteria are not met:
- Write a new continuation that:
- Directly follows from the last sentence of "Prefix Text"
- Matches the style and tone of "Perturbed Generation"
- Maintains factual consistency with "Prefix Text"
OUTPUT ONLY the continuation text, with no explanations or additional text.
——"Prefix Text":
——"Perturbed Generation":
——"Your Continuation":
5.4 Results
Figure 4: Privacy Leakage Analysis with Varying Privacy Budget ($\epsilon$). The figure illustrates the relationship between the privacy budget $\epsilon$ and two key metrics: recovery extent and privacy leakage. The top plot displays the recovery extents of the random document $R(\breve{P})$ and the perturbed document $R(\widetilde{P})$, while the bottom plot shows the privacy leakage, both as functions of the privacy budget $\epsilon$.
Figure 5: Privacy and Utility Tradeoff. This figure illustrates the relationship between nine utility loss metrics (vertical axis) and one privacy leakage metric (horizontal axis). Each subplot shows the results for 24 different privacy budgets $\epsilon$, reflecting the corresponding utility loss and privacy leakage. As $\epsilon$ decreases, privacy leakage increases while utility loss decreases. This clearly demonstrates the tradeoff between privacy leakage and utility loss.
According to our definition of privacy leakage Definition3.1, a recovery extent $R$ closer to 0 indicates a higher amount of recovered private information. As shown in figure 4, $R(\widetilde{P})$ increases as the privacy budget $\epsilon$ increases, suggesting that a higher $\epsilon$ reduces the difficulty of recovering protected prompts, allowing the remote black-box LLM to retrieve more private information from them. In contrast, the $R(\breve{P})$ rate remains relatively stable across different $\epsilon$ values, indicating that random guessing does not affect the amount of recovered private information with changes in the privacy budget. This stability serves as a baseline, highlighting the sensitivity of protected prompt recovery to changes in the privacy budget $\epsilon$ in remote black-box LLMs.
Additionally, privacy leakage ($\epsilon_{p}$) increases as $\epsilon$ rises, consistent with the expectation that a higher privacy budget allows for greater information leakage. At lower values of $\epsilon$, privacy leakage is minimal, and in some cases even negative, indicating strong privacy protection, where the remote black-box LLM’s ability to recover protected prompts is worse than random guessing. However, as $\epsilon$ approaches higher values, privacy leakage increases, suggesting that with a larger privacy budget, the remote black-box LLM is able to recover more private information from the protected prompts, thereby increasing the privacy risk.
Although this result is consistent with our defined relationship between privacy leakage and $\epsilon$, it is important to note that we did not use a specialized recovery algorithm for the experiment. Instead, recovery was performed by sending instructions to the remote black-box LLM. We believe that employing a dedicated iterative recovery algorithm could further improve the accuracy of privacy leakage detection.
Moreover, as shown in Figure 5, this experimental result illustrates the relationship between privacy leakage and utility loss, providing a clear visualization of the trade-off between the two. The figure reveals that as privacy leakage increases—indicating a decrease in the level of privacy protection—all utility metrics (e.g., BERTScore and semantic similarity) exhibit a declining trend. This demonstrates that InferDPT requires a balance between privacy leakage and utility loss: while higher levels of privacy protection effectively reduce privacy leakage, they come at the cost of significant utility loss. Conversely, lower levels of privacy protection improve utility but compromise privacy. Only at an optimal level of privacy protection can both privacy and utility be reasonably balanced. These findings strongly validate our theory that there is an inherent trade-off between privacy leakage and utility loss, aligning well with the "no free lunch" theorem. This discovery offers valuable insights for optimizing privacy-preserving strategies in LLM inference.
@@ 6 Conclusion
To answer this question, we propose a privacy-preserving LLM inference framework. Within this framework, we propose and experimentally verified a No-Free-Lunch Theorem for privacy-preserving LLM inference, demonstrating that the weighted summation of these factors exceeds a constant which is contingent on specific problem and non-zero. This highlights inevitable loss of utility when leakage budget of privacy is excessively constrained.
@@ 7 ACKNOWLEDGMENTS
We thank Yulin Fei for helpful comments. This work was supported by the National Science and Technology Major Project under Grant 2022ZD0115301.
@@
Lemma A.5 (Theoretical Relationship between Recovery Extent and Distortion Extent).
@@
Let $\epsilon_{p}$ be defined in Definition 3.1. Let $\widetilde{P}$ and $\breve{P}$ denote the distorted prompt embedding distribution and the independent random prompt embedding distribution. Then we have,